OpenStudy (big_mensa961):

What is wrong with my work? please,check Use variation parameters to find particular solution for x"+x=cost, x1=cost, x2 =sint

OpenStudy (loser66):

Thanks @Big_Mensa961 but I got it now.

OpenStudy (dinamix):

\[x"+x=0 => r^2+1=0 => r^2=-1 =>r^2=\pm i\]

OpenStudy (radar):

If \[r ^{2}=-1\]

OpenStudy (radar):

How does \[r ^{2}=\pm i\] also ??

OpenStudy (dinamix):

ops i made mistake \[r=\pm i \] ty @radar

OpenStudy (dinamix):

the complete integral of corresponding equation homogeneus x''+x =0 is : \[X=acost+bsint\] and the particular solution the nonhomogneus and since the i is a simple root of the auxiliary equation i seek the paritcular solution in the form \[x* =t(Acost+Bsint)\] and put this solution and romplace it in this \[x''+x=cost\] and find A and B (A not equal a not same that also for b) the general solution is : \[x=X+x*\]

OpenStudy (dinamix):

cuz i found A=0 and B=1/2 when i remplace it @wio could u check the my solution for this equation \[x"+x=cost\] i solve it look up

OpenStudy (dinamix):

\[x= acost+bsint+\frac{ 1 }{ 2 }tsint\]

OpenStudy (dinamix):

@IrishBoy123 could u see my solution for him

OpenStudy (irishboy123):

that seems to be the right answer :-)) https://www.wolframalpha.com/input/?i=x%22%2Bx%3Dcost